If f(x)=5x-1 , find the following
f(a-2)
f(a-2)=

Answer by praseenakos@yahoo.com(507)

QUESTION:

If f(x)=5x-1 , find the following
f(a-2)



ANSWER:

If f(x)=5x-1 , find f(a-2


To find f(a-2), replace x by (a-2) in the given function.

So f(a-2)= 5(a-2) -1

Now multiply 5 with both the terms inside the parenthesis.


==> f(a-2) = 5a - 5*2 -1


==> = 5a - 10 -1


==> = 5a -11 (here -10-1 = -11)


So f(a-2) = 5a -11



Hope you understood.



Regards.


Praseena.